# On the maximal correlation of some stochastic processes

### v2

@@ 1 Introduction
Consider independent continuous random variables $X_{1},X_{2},\ldots,X_{n}$. Let $S\subset T$ be two nested non-empty random subsets of $[n]=\{1,2,\ldots,n\}$, which are independent from $X=(X_{1},X_{2},\ldots,X_{n})$. Let $R$ be the maximal correlation coefficient between $(S,X_{S})$ and $(T,X_{T})$ given by (10), where $X_{S}=(X_{i})_{i\in S}$ and $X_{T}=(X_{j})_{j\in T}$. Denote by $I(Z)=\int_{-\infty}^{\infty}\frac{(f^{\prime}(z))^{2}}{f(z)}\,\mathrm{d}z$ the Fisher information of the continuous random variable $Z$ with the density $f(z)$. Then, we have that
@@
Organization of the paper: In Section 2, we present several useful properties of the maximal correlation coefficients, which are used in the following proofs. In Section 3, we consider maximal correlation coefficients for random walks. We prove Theorem 1.1 and give new proof of Lancaster’s classical result (3) in separate subsections. In Section 4, we consider the maximal correlation coefficients for two dimensional Lévy processes, two dimensional stable processes and two dimensional stable distributions. We first give the proof of Theorem 1.2. Then, we use Theorem 1.2 to prove Theorem 1.3. As applications, we also discuss several examples. These are presented in different subsections. In Section 5, we consider Madiman-Barron type equalities and prove Proposition 1.4, Theorem 1.6, Theorem 1.7, Corollary 1.8 and Theorem 1.9 in separate subsections. Among these results, Proposition 1.4 connects Section 3 and Section 5. More precisely, it can be viewed as an application of Theorem 1.1 and it is a special case of Theorem 1.5. The result of the special case in Proposition 1.4 is actually the motivation for our alternative expression for Madiman-Barron equality. (The original description of Madiman and Barron is a bit different.) We also discover a Dembo-Kagan-Shepp-Yu-type upper bound in Proposition 5.5. There, we replace the summation in (5) by minimum in (55). It implies the upper bound in [BS25, Corollary 2.2] for bivariate Marshall Olkin exponential distribution. Finally, we present two open problems in Section 6.
@@ 2.1 Basic properties
We collect several basic properties in the following:
@@
    In general, the supremum in the definition of the maximal correlation coefficient cannot be replaced by the maximum. For instance, suppose that $M$ and $N$ are independent Poisson random variables of parameter $1$. Let $X=M-N$ and $Y=M-2N$. Then, we have that $R(X,Y)=1$. Indeed, $\lim_{m\to\infty}\rho(1_{X=m},1_{Y=m})=1$. However, for any pair of measurable functions $\varphi$ and $\psi$ such that $E\varphi(X)=E\psi(Y)=0$ and $\operatorname{Var}\varphi(X)=\operatorname{Var}\psi(Y)=1$, we have that $\rho(\varphi(X),\psi(Y))<1$. Rényi [R5́9, Theorem 1, Theorem 2] provided sufficient condition under which the maximal correlation could be attained at some $\varphi(X)$ and $\psi(Y)$.
@@ 2.2 Submultiplicative property
Suppose that $X$ and $Z$ are conditionally independent given $Y$, then
@@
This is the first part of [Yu08, Lemma 2.1]. For self-containedness, we provide a short proof. The result comes from the fact that the operator norm of the composition of two operators is not greater than the product of the operator norms of these two operators. In the proof, the Markov property of $(X,Y,Z)$ is crucially used.
Proof.
[…]
where $\pi_{1},\pi_{2},\pi_{3},\pi_{4}$ and $\pi_{5}$ are orthogonal projections and $\iota$ is the injection. Equivalently, $\pi_{1}:W\mapsto E(W|Z)$, $\pi_{2}:W\mapsto E(W|Y)$, $\pi_{3}:W\mapsto E(W|Y,Z)$, $\pi_{4}:W\mapsto E(W|Z)$ and $\pi_{5}:W\mapsto E(W|Z)$ are the operators of conditional expectations. Hence, we have that
where |||||$\cdot$||| | ⋅ | | is the operator norm. ∎
@@ 2.3 Csáki-Fischer identity
The following result is known as Csáki-Fischer identity [CF63, Theorem 6.2], see also [Wit75, Theorem 1].
@@
Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are independent. Then, we have that
@@
where the conditional covariance $\mathrm{Cov}(U,V|\mathcal{F})$ of two random variables $U,V$ given the sigma field $\mathcal{F}$ is defined by
@@
Because $X_{2}$ and $Y_{1}$ are conditionally independent given $X_{1}$, $E[f(X_{1},X_{2})|\mathcal{F}]$ is a function of $X_{1}$. Similarly, $E[g(Y_{1},Y_{2})|\mathcal{F}]$ is a function of $Y_{1}$. Then, by definition of maximal correlation coefficients, we have that
$|I_{2}|\leq R(X_{1},Y_{1})\sqrt{\operatorname{Var}(E[f(X_{1},X_{2}))|\mathcal{F}])\operatorname{Var}(E[g(Y_{1},Y_{2})|\mathcal{F}])}.$
Next, we look for the upper bound of $I_{1}$. By definition of maximal correlation coefficients and independence between $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$, we have that
$|\operatorname{Cov}(f(X_{1},X_{2}),g(Y_{1},Y_{2})|\mathcal{F})|\leq R(X_{2},Y_{2})\sqrt{\operatorname{Var}(f(X_{1},X_{2})|\mathcal{F})\operatorname{Var}(g(Y_{1},Y_{2})|\mathcal{F})}.$
By taking the expectation on both sides and using Hölder’s inequality, we obtain that
@@
$B_{f}^{X}=\operatorname{Var}(E[f(X_{1},X_{2}))|\mathcal{F}]),B_{g}^{Y}=\operatorname{Var}(E[g(Y_{1},Y_{2})|\mathcal{F}]).$
@@
On the other hand, clearly, we have that
@@
Then, we have the joint distribution of $X_{2},Y_{2}$, see Table 1:
$X_{2}\backslash Y_{2}$1 1
@@
Immediately, we see that $X_{2}$ is independent of $Y_{2}$. Hence, we have that $R(X_{2},Y_{2})=0$. Note that $X_{1}X_{2}=Y_{1}Y_{2}=\sigma X_{1}Y_{1}$ and they are non-degenerate. Consequently, we have that
@@ 2.4 Lower semi-continuity
The maximal correlation coefficient $R(X,Y)$ is fully determined by the joint distribution $\mu$ of $(X,Y)$. Therefore, for a joint distribution $\mu$ of $(X,Y)$, its maximal correlation coefficient $R(\mu)$ is well-defined by $R(\mu)=R(X,Y)$. Suppose that $(X,Y)$ take values in a complete separable metric space $S$ (i.e. Polish space) with the Borel-$\sigma$-field $\mathcal{S}$. Let $\mathcal{P}$ be the space of probability measures on $(S,\mathcal{S})$. We equip $\mathcal{P}$ with the Lévy-Prokhorov metric $\pi$, see [Bil99, Eq. (6.10)] for the definition of this metric. By [Bil99, Theorem 6.8], $(\mathcal{P},\pi)$ is Polish and weak convergence is equivalent to the convergence with respect to the metric $\pi$. In this section, we will show the lower semi-continuity of $\mu\mapsto R(\mu)$ with respect to the metric $\pi$. It is a generalization of [BDK05, Theorem 1].
@@
Suppose that $(X,Y)$ takes values in a Polish space $(S,\mathcal{S})$ and let $\mu$ be the joint distribution of $(X,Y)$. As a function on the Polish space $(\mathcal{P},\pi)$, $\mu\mapsto R(\mu)=R(X,Y)$ is lower semi-continuous.
To prove Lemma 2.3, we need to use the classical result that $L^{p}$ functions can be approximated by bounded continuous functions. We give the precise statement in the following:
Lemma 2.4.
Consider a metric space S with Borel $\sigma$-field $\mathcal{S}$, a bounded measure $\mu$ on $(S,\mathcal{S})$, and a constant $p>0$. Then the bounded continuous functions on S are dense in $L^{p}(S,\mathcal{S},\mu)$. Thus, for any $f\in L^{p},$ there exist some bounded, continuous functions $f_{1},f_{2},\ldots:S\to\mathbb{R}$ with $\|f_{n}-f\|_{p}\to 0$.
The above lemma is precisely Lemma 1.37 in [Kal21].
Proof of Lemma 2.3.
Let $(X_{n},Y_{n})$ be a sequence of random variables with values in the measurable space $(S,\mathcal{S})$. Suppose that $(X_{n},Y_{n})$ converges weakly to $(X,Y)$ as $n\to+\infty$. It suffices to show that
$R(X,Y)\leq\liminf_{n\to+\infty}R(X_{n},Y_{n}).$
For any $\varepsilon>0$, by definition of the maximal correlation coefficient, there exists $\varphi$ and $\psi$ such that
$R(X,Y)\leq\rho(\varphi(X),\psi(Y))+\varepsilon$
and that $\operatorname{Var}(\varphi(X))>0$ and $\operatorname{Var}(\psi(Y))>0$. By Lemma 2.4 for $p=2$, for any $\delta>0$, there exist bounded continuous functions $f$ and $g$ such that
$E[(f(X)-\varphi(X))^{2}]<\delta,E[(g(Y)-\psi(Y))^{2}]<\delta.$
$\lim_{n\to\infty}\rho(f(X_{n}),g(Y_{n}))=\rho(f(X),g(Y)).$
By definition of maximal correlation coefficients,
$\rho(f(X_{n}),g(Y_{n}))\leq R(X_{n},Y_{n}).$
Hence, for any $\varepsilon>0$, we have that
By taking $\varepsilon\to 0$, we obtain the desired result. ∎
@@
Later after we independently prove Lemma 2.3, we find a very similar statement in [KA16, Section II-A]. The idea of proofs are similar. However, there are several restrictions in their statement and proof. Firstly, they require that $X$ and $Y$ take values in finite state spaces. Secondly, the convergence of $(X_{n},Y_{n})$ towards $(X,Y)$ is assumed to be the convergence with respect to total variation distance instead of weak convergence.
@@ 3 Random walks
In this subsection, we prove Theorem 1.1. As we have explained, the first equality is a direct consequence of Csáki-Fischer identity, see Theorem 2.2. To prove the second equality, for natural numbers $m\geq 1$, we define continuous time processes
@@
Then, we have that
@@
As $m\to\infty$, the process $(S^{(m)},T^{(m)})$ converges to $(S,T)$ in the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$ by [Bil99, Theorem 16.1], where $D_{\mathbb{R}^{2}}[0,\infty)$ is the space of càdlàg functions $\omega:[0,\infty)\to\mathbb{R}^{2}$ endowed with the Skorokhod topology. Hence, by Lemma 2.3, we have that
@@
3.2 New proof of (3)
In this subsection, we give a new proof of (3) as a consequence of Theorem 1.1, the central limit theorem and Lemma 2.3.
Fix a Gaussian vector $(X,Y)$ such that the covariance is $\Sigma$ and the Pearson correlation coefficient is $r$.
@@ 3.2 New proof of (3)
Then, for a non-constant function $f$ on $\{a,b\}$ and a non-constant function $g$ on $\{c,d\}$, the Pearson correlation coefficient between $f(\xi)$ and $g(\eta)$ is
$\operatorname{sgn}(f(a)-f(b))\operatorname{sgn}(g(c)-g(d))\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}/\sqrt{p_{a}p_{b}p_{c}p_{d}}.$
Then, we have that $R(\xi,\eta)=\left|\rho(\xi,\eta)\right|=\left.\left|\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}\right.\right|/\sqrt{p_{a}p_{b}p_{c}p_{d}}$. By proper choices of $a,b,c,d$ and the probability mass function $p$, we may assume that the covariance matrix of $(\xi,\eta)$ is precisely $\Sigma$, and the Pearson correlation coefficient equals to $r$.
Let $(\xi_{n},\eta_{n})_{n\geq 1}$ be i.i.d. random vectors such that $(\xi_{n},\eta_{n})$ has the same distribution as $(\xi,\eta)$. Define a two-dimensional random walk $(S,T)$ via (6). Then, by Theorem 1.1, we have that
@@
By the central limit theorem, as $m\to\infty$, $\left(\frac{S_{m}-ES_{m}}{\sqrt{m}},\frac{T_{m}-ET_{m}}{\sqrt{m}}\right)$ converges to $(X,Y)$ in distribution. By Lemma 2.3, we see that
@@
On the other hand, by definition, we have that $R(X,Y)\geq|\rho(X,Y)|$. Thus, (3) is proved.
4 Lévy processes and stable distributions

### v3

@@ 1.3 Extensions of the Dembo-Kagan-Shepp-Yu inequality
Let $X_{1},X_{2},\ldots,X_{n}$ be independent real-valued continuous random variables. Let $S\subset T$ be two nested non-empty random subsets of $n]=\{1,2,\ldots,n\}$, which are independent of $X=(X_{1},X_{2},\ldots,X_{n})$. Let $R$ be the maximal correlation coefficient between $(S,X_{S})$ and $(T,X_{T})$ given by (8), where $X_{S}=(X_{i})_{i\in S}$ and $X_{T}=(X_{j})_{j\in T}$. Denote by $I(Z)=\int_{-\infty}^{\infty}\frac{(f^{\prime}(z))^{2}}{f(z)}\,\mathrm{d}z$ the Fisher information of the continuous random variable $Z$ with the density $f(z)$. Then we have that
@@
Organization of the paper: In Section 2, we present several useful properties of maximal correlation coefficients. The proofs of Lemma 2.1 and Lemma 2.3 are postponed to the appendix. In Section 3, we prove Theorem 1.1 and give a new proof of Lancaster’s classical result (3). In Section 4, we first prove Theorem 1.2, and then use Theorem 1.2 to prove Theorem 1.3. We also give several examples in this section. In Section 5, we prove the lower bounds of Theorem 1.4, Theorem 1.5, Theorem 1.6, Corollary 1.7 and Theorem 1.8 in separate subsections. We also give a Dembo-Kagan-Shepp-Yu-type upper bound in Proposition 5.5, where the summation in (7) is replaced by the minimum in (55). Finally, we present two open problems in Section 6.
@@ 2.1 Basic properties
We collect several basic properties as follows:
@@
    In general, the supremum in the definition of the maximal correlation coefficient cannot be replaced by the maximum. For example, suppose that $M$ and $N$ are independent Poisson random variables of parameter $1$. Let X=MNX=M-N and Y=M2NY=M-2N. Then we have that $R(X,Y)=1$. Indeed, $\lim_{m\to+\infty}\rho(1_{X=m},1_{Y=m})=1$. However, for any measurable functions $\varphi$ and $\psi$ satisfying $E[\varphi(X)]=E[\psi(Y)]=0$ and $\operatorname{Var}(\varphi(X))=\operatorname{Var}(\psi(Y))=1$, we have that $\rho(\varphi(X),\psi(Y))<1$. Rényi [R5́9, Theorems 1 and 2] established sufficient conditions for the supremum to be attained.
@@ 2.2 Submultiplicative property
Suppose that $X$ and $Z$ are conditionally independent given $Y$. Then
@@
This is the first part of [Yu08, Lemma 2.1]. For self-containedness, we provide a short proof in Appendix A.
@@ 2.3 Csáki-Fischer identity
The following result is known as the Csáki-Fischer identity [CF63, Theorem 6.2], see also [Wit75, Theorem 1].
@@
Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are independent. Then we have that
@@
where the conditional covariance $\mathrm{Cov}(U,V|\mathcal{F})$ of two random variables $U,V$ given the sigma-field $\mathcal{F}$ is defined by
@@
Due to the conditional independence of $X_{2}$ and $Y_{1}$ given $X_{1}$, $E[f(X_{1},X_{2})|\mathcal{F}]$ is a function of $X_{1}$. Similarly, $E[g(Y_{1},Y_{2})|\mathcal{F}]$ is a function of $Y_{1}$. By the definition of maximal correlation coefficients, we have that
$|I_{2}|\leq R(X_{1},Y_{1})\sqrt{\operatorname{Var}(E[f(X_{1},X_{2})|\mathcal{F}])\operatorname{Var}(E[g(Y_{1},Y_{2})|\mathcal{F}])}.$
Next, we look for the upper bound of $I_{1}$. By the definition of maximal correlation coefficients and independence between $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$, we have that
[…]
By taking the expectation on both sides and using the Cauchy-Schwarz inequality, we obtain that
@@
$B_{f}^{X}=\operatorname{Var}(E[f(X_{1},X_{2})|\mathcal{F}]),B_{g}^{Y}=\operatorname{Var}(E[g(Y_{1},Y_{2})|\mathcal{F}]).$
@@
On the other hand, we have that
@@
Then the joint distribution of $X_{2},Y_{2}$ is given in Table 1.
[…]
@@
Immediately, we see that $X_{2}$ is independent of $Y_{2}$. Hence, we have that $R(X_{2},Y_{2})=0$. Note that $X_{1}X_{2}=Y_{1}Y_{2}=\sigma X_{1}Y_{1}$, and they are non-degenerate. Consequently, we have that
@@ 2.4 Lower semi-continuity
The maximal correlation coefficient $R(X,Y)$ is fully determined by the joint distribution $\mu$ of $(X,Y)$. Therefore, for a joint distribution $\mu$ of $(X,Y)$, its maximal correlation coefficient $R(\mu)$ is well-defined by $R(\mu)=R(X,Y)$. Suppose that $(X,Y)$ takes values in a complete separable metric space $S$ (i.e., Polish space) with the Borel $\sigma$-field $\mathcal{S}$. Let $\mathcal{P}$ be the space of probability measures on $(S,\mathcal{S})$. We equip $\mathcal{P}$ with the Lévy-Prokhorov metric $\pi$, see [Bil99, Eq. (6.10)] for the definition of this metric. By [Bil99, Theorem 6.8], $(\mathcal{P},\pi)$ is Polish, and weak convergence is equivalent to convergence with respect to the metric $\pi$. In this section, we will show the lower semicontinuity of $\mu\mapsto R(\mu)$ with respect to the metric $\pi$. It is a generalization of [BDK05, Theorem 1].
@@
Suppose that $(X,Y)$ takes values in a Polish space $(S,\mathcal{S})$ and let $\mu$ be the joint distribution of $(X,Y)$. As a function on the Polish space $(\mathcal{P},\pi)$, $\mu\mapsto R(\mu)=R(X,Y)$ is lower semicontinuous.
As noted by an anonymous referee, the semicontinuity is quite natural since the supremum operator preserves semicontinuity. However, for self-containedness, we provide a proof in Appendix B.
@@
After proving Lemma 2.3, we became aware of a very similar statement in [KA16, Section II-A]. While the proof ideas are similar, their result has restrictions. Firstly, $X$ and $Y$ are restricted to finite state spaces. Secondly, the convergence of $(X_{n},Y_{n})$ towards $(X,Y)$ is assumed to be the convergence with respect to the total variation distance instead of the weak convergence.
@@ 3 Random walks
In this subsection, we prove Theorem 1.1. As we have explained, the first equality is a direct consequence of the Csáki-Fischer identity, see Theorem 2.2. To prove the second equality, for natural numbers $m\geq 1$, we define two continuous-time processes
@@
Then we have that
@@
As $m\to\infty$, the process $(S^{(m)},T^{(m)})$ converges to $(S,T)$ in the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$ by [JS03, Theorem 1.14], where $D_{\mathbb{R}^{2}}[0,\infty)$ denotes the space of càdlàg functions $\omega:[0,\infty)\to\mathbb{R}^{2}$ endowed with the Skorokhod topology. Therefore, by Lemma 2.3, we have that
@@
3.2 A new proof of (3)
In this subsection, we provide a new proof of (3) based on Theorem 1.1, the central limit theorem and Lemma 2.3.
Let $(X,Y)$ be a two-dimensional Gaussian vector with covariance matrix $\Sigma$ and Pearson correlation coefficient $r$.
@@ 3.2 A new proof of (3)
Then for a non-constant function $f$ on $\{a,b\}$ and a non-constant function $g$ on $\{c,d\}$, the Pearson correlation coefficient between $f(\xi)$ and $g(\eta)$ is given by
$\operatorname{sgn}(f(a)-f(b))\operatorname{sgn}(g(c)-g(d))\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}/\sqrt{p_{a}p_{b}p_{c}p_{d}},$
where $p_{a}=P(\xi=a)$, $p_{b}=P(\xi=b)$, $p_{c}=P(\eta=c)$ and $p_{d}=P(\eta=d)$. Consequently,
$R(\xi,\eta)=\left|\rho(\xi,\eta)\right|=\left.\left|\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}\right.\right|/\sqrt{p_{a}p_{b}p_{c}p_{d}}.$ (12)
By appropriately choosing $a$, $b$, $c$, $d$ and the probability mass function $p$, we can ensure that $(\xi,\eta)$ has covariance matrix $\Sigma$ and Pearson correlation coefficient $r$.
Let $(\xi_{n},\eta_{n})_{n\geq 1}$ be i.i.d. random vectors such that $(\xi_{n},\eta_{n})$ has the same distribution as $(\xi,\eta)$. Define a two-dimensional random walk $(S,T)$ as in (6). By Theorem 1.1, we have that
@@
By the central limit theorem, $\left(\frac{S_{m}-ES_{m}}{\sqrt{m}},\frac{T_{m}-ET_{m}}{\sqrt{m}}\right)$ converges to $(X,Y)$ in distribution as $m\to\infty$. By Lemma 2.3, we see that
@@
On the other hand, by definition, we have that $R(X,Y)\geq|\rho(X,Y)|$. Therefore, (3) holds.
4 Two-dimensional Lévy processes and bivariate stable distributions
Recall the Lévy-Khinchine formula for a two-dimensional Lévy process $(X_{t},Y_{t})_{t\geq 0}$:
[…]
$+\left.\left.\int_{\mathbb{R}^{2}\setminus\{(0,0)\}}\left(e^{i(u_{1}x+u_{2}y)}-1-i(u_{1}x+u_{2}y)I_{(0,1]}(x^{2}+y^{2})\right)\,\nu(\mathrm{d}x,\mathrm{d}y)\right\}\right).$ (13)
Here, $\Sigma=(\Sigma_{jk})_{j,k=1,2}$ is the covariance matrix of the Brownian motion in the Lévy-Itô decomposition, and $\nu(\mathrm{d}x,\mathrm{d}y)$ is the Lévy measure on $\mathbb{R}^{2}\setminus\{(0,0)\}$ satisfying
$\int_{\mathbb{R}^{2}\setminus\{(0,0)\}}\min(1,x^{2}+y^{2})\,\nu(\mathrm{d}x,\mathrm{d}y)<\infty.$
A similar formula holds for an infinitely divisible random vector. For a two-dimensional stable random vector, the Lévy measure $\nu$ takes the form
$\nu(B)=\int_{0}^{\infty}\int_{\mathbb{R}/(2\pi\mathbb{Z})}I_{B}(r\cos\theta,r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta),\quad B\in\mathcal{B}(\mathbb{R}^{2}\setminus\{(0,0)\}),$ (14)
where the spectral measure $\tau(\mathrm{d}\theta)$ is a finite Borel measure and $\mathcal{B}(\mathbb{R}^{2}\setminus\{(0,0)\})$ denotes the Borel $\sigma$-field of $\mathbb{R}^{2}\setminus\{(0,0)\}$.